# Quantum LDPC codes for erasure-biased atomic quantum processors

### v1

@@ Abstract
Identifying the best families of quantum error correction (QEC) codes for near-term experiments is key to enabling fault-tolerant quantum computing. Ideally, such codes should have low overhead in qubit number, high physical error thresholds, and moderate requirements on qubit connectivity to simplify experiments, while allowing for high logical error suppression. Quantum Low-Density Parity-Check (LDPC) codes have been recently shown to provide a path towards QEC with low qubit overhead and small logical error probabilities. Here, we demonstrate that when the dominant errors are erasures – as can be engineered in different quantum computing architectures – quantum LDPC codes additionally provide high thresholds and even stronger logical error suppression in parameter regimes that are accessible to current experiments. Using large-scale circuit-level QEC simulations, we benchmark the performance of two families of high-rate quantum LDPC codes, namely Clifford-deformed La-cross codes and Bivariate Bicycle codes, under a noise model strongly biased towards erasure errors. Both codes outperform the surface code by offering up to orders of magnitude lower logical error probabilities. Interestingly, we find that this decrease in the logical error probability may not be accompanied by an increase in the code threshold, as different QEC codes benefit differently from large erasure fractions. While here we focus on neutral atom qubits, the results also hold for other quantum platforms, such as trapped ions and superconducting qubits, for which erasure conversion has been demonstrated.
@@ I Introduction
Bivariate Bicycle [9] and La-cross [11] quantum LDPC codes are candidates for near-term experimental realizations in neutral atom [11, 13, 12], trapped-ion [17] and superconducting [9, 12] platforms. They have been shown to offer significantly lower overhead and better error correction performance than the surface code [9, 11], while only requiring few non-local interactions of moderate extent. Still, the implementation of high-rate quantum LDPC codes remains challenging as these codes typically suffer from lower circuit-level thresholds than surface and color codes, as their higher stabilizer weight increases the number of error locations.
@@
The XZZX surface code has also been shown to offer significant enhancements in QEC performance when a large fraction of errors are erasures [26, 27], i.e. heralded qubit losses [28]. Erasures are less detrimental than generic depolarizing errors, since their actual location in the quantum register is known. The desirable scenario where qubits are strongly biased towards erasure errors can be experimentally realized via erasure conversion mechanisms [26, 22, 29] to reveal the locations of losses. Examples of these mechanisms have been originally introduced to mitigate the effect of Rydberg leakages in Alkaline-earth(-like) atom qubits, such as Yb171 [26, 27, 30], and atom losses [22, 29, 31, 32]. Beyond neutral atom platforms, erasure conversion mechanisms have also been demonstrated for trapped ions [33] and superconducting qubits [34].
Identifying the best QEC code family that can be implemented in near-term quantum processors is currently a key open challenge. Ideally, such a code family should have low overhead, high circuit-level threshold, and moderately long-range connectivity, while allowing for high logical error suppression. In this work, we find that Clifford-deformed La-cross codes and, partially, Bivariate Bicycle codes fulfill all these requirements under a noise model strongly biased towards erasure errors. We use extensive QEC circuit-level numerical simulations to show that there exists an experimentally achievable range of physical error probabilities where both codes outperform the surface code in any respect, offering up to orders of magnitude lower logical error probabilities. In particular, Clifford-deformed La-cross codes offer thresholds that are comparable to those of the surface code under both unbiased and biased erasures, while Bivariate Bicycle codes do not display significant threshold improvement for any finite fraction of unbiased erasures. The noise model that we consider in our work can be experimentally realized in Alkaline-earth(-like) atom qubits [26].
@@ II Overview of erasure conversion protocols for neutral atoms
In neutral atom quantum processors [35, 36, 37, 38, 39], atoms – generally Alkali or Alkaline-earth(-like) – are loaded from a magneto-optical trap into a optical tweezer array generated by a spatial light modulator (SLM) and rearranged in a defect-free configuration via 2D acousto-optic deflectors (AODs) [40]. Single-qubit rotations are performed via robust Raman laser-excitations with more than $99.95\%$ fidelity [40]. Let atoms be three-level systems with computational states $\{|0\rangle,|1\rangle\}$ and auxiliary Rydberg state $|r\rangle$. Controlled-Z (CZ) entangling gates are usually realized by shining global laser pulses coupling the $|1\rangle\leftrightarrow|r\rangle$ states of each atom, such that the target atom in state $|1\rangle$ acquires a $\pi$-phase conditionally on the control atom also being in state $|1\rangle$ [41]. The strong Van der Waals interaction between the two atoms naturally prevents them to be simultaneously Rydberg-excited [36]. Recently, CZ gate fidelities in neutral atom qubits have recorded values of $99.4-99.7\%$ [42, 43, 44].
A particularly detrimental error source affecting neutral atom experiments is qubit loss. Several mechanisms can cause atoms to be lost during the computation, such as background gas collisions, heating by repeated gate execution or atom shuttling, and atom drift by imperfect trapping or imperfect recapture [22, 29]. That is because, usually, optical traps are trapping when atoms are in their ground state and anti-trapping when they are in a Rydberg-excited state. Consequently, they have to be turned off during the gate execution and then turned on again, with a non-zero probability that the atom is lost during the process. Two other error sources causing qubit loss are Rydberg leakages and hyperfine leakages [29]. The former refers to atoms unintentionally left in some Rydberg state after the gate execution, either by over-rotation or by black-body induced transitions to neighboring Rydberg states. The latter refers to decays to hyperfine Zeeman sublevels. In both cases, atoms leave the computational subspace, resulting in information loss.
@@ III.1 La-cross codes
where now $(H_{Q})_{\alpha\beta}=1$ iff the $\alpha$th stabilizer acts non-trivially on the $\beta$th qubit. In Eq. 1, the left part of $H_{Q}$ describes $X$-stabilizers and the right one $Z$-stabilizers. Quantum parameters read $N=n_{1}n_{2}+r_{1}r_{2}$, $K=k_{1}k_{2}+k_{1}^{T}k_{2}^{T}$, $D\geq\min\{d_{1},d_{2},d_{1}^{T},d_{2}^{T}\}$ [45] and the number of $X$- and $Z$-stabilizers is equal to $n_{1}r_{2}$ and $n_{2}r_{1}$, respectively. The resulting code is of Calderbank-Shor-Steane (CSS) type [47, 48], which means that the stabilizers of the code are products of only $X$ or only $Z$ Pauli operators. We can now perform a Clifford deformation consisting of applying Hadamard rotations to alternating data qubits [23]. Upon this transformation, the above parity-check matrix becomes [24]
@@
La-cross quantum LDPC codes are HGP codes built from equal cyclic classical codes with a generating polynomial $h(x)=1+x+x^{k}$, Hamming distance $d$, and classical parity-check matrix $H=\text{circ}(1,1,0,\dots,0,1,0,\dots,0)\in\mathbb{F}_{2}^{(n-k)\times n}$ (with the locations of the non-zero entries dictated by the generating polynomial, i.e. $0,1,k$) which is circulant rectangular [11]. La-cross codes have code parameters $N=(n-k)^{2}+n^{2}$, $K=k^{2}$, $D=d$ and weight-$6$ stabilizers consisting of four nearest-neighbor and two symmetrically separated distant qubits – or stabilizer legs. The length of the stabilizer legs is equal to $k$. Classical seeds having circulant rectangular parity-check matrix guarantee the resulting La-cross codes to enjoy open boundary conditions [11]. Since for La-cross codes the classical seed codes are identical, we can drop the indices and use the classical parameter $k$ as a label for different code instances. In fact, for equal rectangular seeds encoding $k$ bits, the resulting HGP code will encode $k^{2}$ logical qubits, hence the parameter $k$ sets both the long-range connectivity and the encoding rate of the resulting code.
@@
Figure 2: Cumulative logical error probability normalized by the number of QEC rounds for $k=2$(a), $k=3$(b) and $k=4$(c) La-cross codes (solid lines) under a fraction $R_{e}=0.98$ of unbiased erasure errors and $R_{p}=1-R_{e}=0.02$ of Pauli errors. A comparison with surface codes under the same noise model and at equal number of physical and logical qubits (dashed lines of the same color and marker style) is also shown. Codes sharing the same number of physical qubits, $N$, and logical qubits, $K$, are denoted with the same color. These results show that La-cross codes have high threshold (black stars) and significantly outperform the surface code in terms of QEC performance below physical error probabilities of $\sim 10^{-2}$, for the code distances we have considered. The insets show how the thresholds of La-cross codes increase by increasing the unbiased erasure fraction for $R_{e}=0.40,0.75,0.90,0.98,1.0$. La-cross code instances shown in these plots are: $[[34,4,3]],[[52,4,4]],[[100,4,5]],[[130,4,6]]$ for $k=2$; $[[45,9,3]],[[65,9,4]],[[149,9,5]],[[225,9,6]]$ for $k=3$; $[[80,16,3]],[[106,16,4]],[[136,16,5]],[[208,16,6]]$ for $k=4$. Error bars on the data are standard deviations associated with the Monte Carlo error correction numerical simulations. For both La-cross and surface code, BP+OSD decoder was used.
@@ III.2 Bivariate Bicycle codes
where $A=A_{1}+A_{2}+A_{3}$ and $B=B_{1}+B_{2}+B_{3}$ are matrix trinomials with $A_{i}$ and $B_{i}$ powers of $x=S_{\ell}\otimes I_{m}$ and $y=I_{m}\otimes S_{\ell}$, respectively. Such a construction ensures the resulting code to be of CSS-type and to have periodic boundary conditions and weight-$6$ stabilizers, as $H_{Q}$ is square and each row contains only six non-zero entries. Bivariate Bicycle code parameters are: $N=2\ell m$, $K=2\times\text{dim}(\text{ker}(A)\cap\text{ker}(B))$, and $D=\text{min}\{|v|,\;v\in\text{ker}([A|B])\,\backslash\,\text{rowspace}([B^{T}|A^{T}])$ which can be computed using the integer linear programming method [9].
@@ III.3 Error correction simulations
In this work we have evaluated the QEC performance of quantum LDPC codes via extensive circuit-level numerical simulations using the Clifford simulator Stim [49]. For Clifford-deformed La-cross codes, open boundary conditions, unrotated array configuration and syndrome extraction scheme depicted in Fig. 1(a) (inset) are assumed [11]. Bivariate Bicycle codes are simulated with periodic boundary conditions, unrotated array configuration and standard syndrome extraction scheme. In both cases, the qubit register is first prepared to state $|0\rangle^{\otimes N}$ and – only for Clifford-deformed La-cross codes – Hadamard (H) rotations are applied on alternating data qubits. We apply CNOT/CZ gates between each data and ancilla qubit and then we measure the ancillas in the proper basis. To ensure fault-tolerance against errors on the ancilla qubits [4], the process is repeated $D$ times for a total of $D$ stabilizer measurement rounds, being $D$ the code distance.
We focus on the effects of two-qubit gate errors, which are usually the most detrimental error mechanism, while we assume single-qubit, state preparation and measurement gates to be perfect, as in Ref. [26]. Idle errors are also neglected. In particular, we assume that each two-qubit gate can experience either a Pauli error with probability $p_{p}=p(1-R_{e})$ or an erasure with probability $p_{e}=pR_{e}$, being $R_{e}$ the fraction of erasure errors. Pauli errors are randomly drawn from $\{I,X,Y,Z\}^{\otimes 2}\backslash\{I\otimes I\}$ with probability $p_{p}/15$, corresponding to the standard two-qubit depolarizing noise. Unbiased erasure errors are instead modeled as follows: When a two-qubit gate is erased, both atoms – data and ancilla – are reset to the maximally mixed state $I/2$, that is we draw errors at random from $\{I,X,Y,Z\}^{\otimes 2}$ with probability $1/16$, conditionally on an erasure to have occurred. Experimentally, this is consistent with replacing the erased atoms with fresh ones initialized to the maximally mixed state from an external reservoir, e.g. via movable tweezers [26]. In addition, reloading both data and ancilla atoms involved in the erased gate prevents correlated errors to spread during the computation.
@@
For both noise models, we assume that mid-circuit ground-state measurements can be interleaved after any gate layer of the syndrome extraction circuit, ensuring that the failed CNOT/CZ gates are heralded along with the error locations and that the erased atoms can be reloaded, as we model in the simulations.
We perform extensive Monte Carlo numerical simulations to probe the code tolerance to errors and use minimum-sum Belief Propagation with Ordered Statistics Decoding (BP+OSD) [50, 51] to process syndrome information of both quantum LDPC codes and surface codes. Although significantly slower than standard matching-based decoders for the surface code, BP+OSD has proven to provide slightly higher threshold and lower logical failure rate [11]. We normalize the data by the number of QEC rounds, $D$, i.e. we plot the cumulative logical failure probability (the probability that any logical qubit fails) as $P_{L}=1-(1-p_{L})^{1/D}$, being $p_{L}$ the ratio between the number of decoder failures and the total number of shots. For surface codes, the logical failure rate is computed as $P^{K}_{L}=1-(1-P_{L})^{K}$, being $K$ the number of logical qubits of the compared quantum LDPC code.
Figure 3: Cumulative logical error probability normalized by the number of QEC rounds for $k=2$(a), $k=3$(b) and $k=4$(c) La-cross code (darker lines) under a fraction $R_{e}=0.98$ of biased erasure errors and $R_{p}=1-R_{e}=0.02$ of Pauli errors. A comparison with La-cross codes under $R_{e}=0.98$ of unbiased erasure errors (lighter lines) is also shown, showing the benefit in the threshold (black stars) and in the logical error probability from a biased and easier-to-decode noise model. La-cross code instances shown in these plots are: $[[34,4,3]],[[52,4,4]],[[100,4,5]],[[130,4,6]]$ for $k=2$; $[[45,9,3]],[[65,9,4]],[[149,9,5]],[[225,9,6]]$ for $k=3$; $[[80,16,3]],[[106,16,4]],[[136,16,5]],[[208,16,6]]$ for $k=4$. Error bars on the data are standard deviations associated with the Monte Carlo QEC simulations. For both La-cross and surface code, BP+OSD decoder was used.
@@
circuit-level, unbiased erasure $1.6\%$ ($1.1\%^{*}$) $5.6\%$ ($4.6\%^{*}$) $6.8\%$ ($5.4\%^{*}$) $0.40\%$ $4.0\%$ $5.6\%$ $0.45\%$ $4.2\%$ $5.9\%$ $0.50\%$ $4.6\%$ $6.2\%$
circuit-level, biased erasure, native gates $1.6\%$ ($1.1\%^{*}$) $10.1\%$ ($8.7\%^{*}$) $11.7\%$ ($10.6\%^{*}$) $0.40\%$ $6.8\%$ $9.1\%$ $0.45\%$ $7.0\%$ $9.4\%$ $0.50\%$ $7.2\%$ $9.6\%$
@@
We show in Fig. 2 results for $k=2$(a), $k=3$(b) and $k=4$(c) Clifford-deformed La-cross codes under a fraction of $R_{e}=0.98$ unbiased erasure errors. We plot their cumulative logical error probability, $P_{L}$, as a function of the physical error probability, $p$, for different code sizes (solid lines). We compare these codes with surface codes of equal/similar number of physical qubits and equal number of logical qubits under the same noise model (dashed lines with the same color and marker style).
@@
It has been shown in Ref. [11] that La-cross codes under depolarizing noise have circuit-level thresholds $p_{th}\approx 0.4-0.5\%$, while for the surface code $p_{th}\approx 1.6\%$. Here, our results show that, under large unbiased erasure fractions, the thresholds of the two code families are comparable. That is, we find thresholds of $p_{th}\approx 4.0-4.6\%$ for La-cross codes [black stars in Fig. 2(a)-(c)] and $p_{th}\approx 5.6\%$ for the surface code. This corresponds to an approximate improvement of ×$\times 10$ against ×$\times 5$, showing that the absolute threshold increase of La-cross codes is larger than that of the surface code.

### v2

# Quantum low-density parity-check codes for erasure-biased atomic quantum processors

@@ Abstract
Identifying the best families of quantum error correction (QEC) codes for near-term experiments is key to enabling fault-tolerant quantum computing. Ideally, such codes should have low overhead in qubit number, high physical error thresholds, and moderate requirements on qubit connectivity to simplify experiments, while allowing for high logical error suppression. Quantum Low-Density Parity-Check (LDPC) codes have been recently shown to provide a path towards QEC with low qubit overhead and small logical error probabilities. Here, we demonstrate that when the dominant errors are erasures – as can be engineered in different quantum computing architectures – quantum LDPC codes additionally provide high thresholds and even stronger logical error suppression in parameter regimes that are accessible to current experiments. Using large-scale QEC numerical simulations, we benchmark the performance of two families of high-rate quantum LDPC codes, namely Clifford-deformed La-cross codes and Bivariate Bicycle codes, under a noise model strongly biased towards erasure errors. Both codes outperform the surface code by offering up to orders of magnitude lower logical error probabilities. Interestingly, we find that this decrease in the logical error probability may not be accompanied by an increase in the code threshold, as different QEC codes benefit differently from large erasure fractions. While here we focus on neutral atom qubits, the results also hold for other quantum platforms, such as trapped ions and superconducting qubits, for which erasure conversion has been demonstrated.
@@ I Introduction
Bivariate Bicycle [9] and La-cross [11] quantum LDPC codes are candidates for near-term experimental realizations in neutral atom [11, 13, 12], trapped-ion [17] and superconducting [9, 12] platforms. They have been shown to offer significantly lower overhead and better error correction performance than the surface code [9, 11], while only requiring few non-local interactions of moderate extent.
Still, the implementation of high-rate quantum LDPC codes remains challenging, as these codes typically suffer from lower circuit-level thresholds than surface and color codes, as their higher stabilizer weight increases the number of error locations.
@@
The XZZX surface code has also been shown to offer significant enhancements in QEC performance when a large fraction of errors are erasures [26, 27], i.e. heralded qubit losses [28]. Erasures are less detrimental than generic depolarizing errors, since their actual location in the quantum register is known. The desirable scenario where qubits are strongly biased towards erasure errors can be experimentally realized via erasure conversion mechanisms [26, 22, 29, 30, 31] to reveal the locations of losses. Examples of these mechanisms have been originally introduced to mitigate the effect of Rydberg leakages in Alkaline-earth(-like) atom qubits, such as Yb171 [26, 27, 32], and atom losses [22, 29, 33, 34]. Beyond neutral atom platforms, erasure conversion mechanisms have also been demonstrated for trapped ions [35, 36, 37] and superconducting qubits [38, 39, 40].
Identifying the best QEC code family that can be implemented in near-term quantum processors is currently a key open challenge. Ideally, such a code family should have low overhead, high circuit-level threshold, and moderately long-range connectivity, while allowing for high logical error suppression. In this work, we move one step forward in this direction by showing that Clifford-deformed La-cross codes and, partially, Bivariate Bicycle codes fulfill all these requirements under a noise model strongly biased towards erasure errors. We consider two different types of erasure errors, namely unbiased erasures and biased erasures [26, 27]. The former refers to the case where the erased qubits are reloaded in the the maximally mixed state, the latter to that where the erased qubits are reloaded in the first excited state, hence introducing – upon Pauli twirling approximation – an additional bias towards Pauli $Z$ errors. We use extensive QEC circuit-level numerical simulations with two-qubit gate errors and show that there exists an experimentally achievable range of physical error probabilities where both codes outperform the surface code in any respect, offering up to orders of magnitude lower logical error probabilities. In particular, Clifford-deformed La-cross codes offer thresholds that are comparable to those of the surface code under both unbiased and biased erasures. Instead, Bivariate Bicycle codes do not display significant threshold improvement for any finite fraction of unbiased erasures, while still offering competitive logical error probabilities below threshold. The noise model that we consider in our work can be experimentally realized in Alkaline-earth(-like) atom qubits [26].
@@ II Overview of erasure conversion protocols for neutral atoms
In neutral atom quantum processors [41, 42, 43, 44, 45], atoms – generally Alkali or Alkaline-earth(-like) – are loaded from a magneto-optical trap into a optical tweezer array generated by a spatial light modulator (SLM) and rearranged in a defect-free configuration via 2D acousto-optic deflectors (AODs) [46]. Single-qubit rotations are performed via robust Raman laser-excitations with more than $99.95\%$ fidelity [46]. Let atoms be three-level systems with computational states $\{|0\rangle,|1\rangle\}$ and auxiliary Rydberg state $|r\rangle$. Controlled-Z (CZ) entangling gates are usually realized by shining global laser pulses coupling the $|1\rangle\leftrightarrow|r\rangle$ states of each atom, such that the target atom in state $|1\rangle$ acquires a $\pi$-phase conditionally on the control atom also being in state $|1\rangle$ [47]. The strong Van der Waals interaction between the two atoms naturally prevents them to be simultaneously Rydberg-excited [42]. Recently, CZ gate fidelities in neutral atom qubits have recorded values of $99.4-99.8\%$ [48, 49, 50, 51].
A particularly detrimental error source affecting neutral atom experiments is qubit loss. Several mechanisms can cause atoms to be lost during the computation, such as background gas collisions, heating by repeated gate execution or atom shuttling, and atom drift by imperfect trapping or imperfect recapture [22, 29]. That is because, usually, optical traps are trapping when atoms are in their ground state and anti-trapping when they are in a Rydberg-excited state. Consequently, the optical traps have to be turned off during the gate execution and then turned on again, with a non-zero probability that the atom is lost during the process. Two other error sources causing qubit loss are Rydberg leakages and hyperfine leakages [29]. The former refers to atoms unintentionally left in some Rydberg state after the gate execution, either by over-rotation or by black-body induced transitions to neighboring Rydberg states. The latter refers to decays to hyperfine Zeeman sublevels. In both cases, atoms leave the computational subspace, resulting in information loss.
@@ III.1 La-cross codes
[…]
@@
La-cross quantum LDPC codes are HGP codes built from equal cyclic classical codes with a generating polynomial $h(x)=1+x+x^{k}$, Hamming distance $d$, and classical parity-check matrix $H=\text{circ}(1,1,0,\dots,0,1,0,\dots,0)\in\mathbb{F}_{2}^{(n-k)\times n}$ (with the locations of the non-zero entries dictated by the generating polynomial, i.e. $0,1,k$) which is circulant rectangular [11]. It has been shown in Ref. [11] that classical codes having circulant rectangular parity-check matrix guarantee the resulting La-cross codes to enjoy open boundary conditions and code parameters $N=(n-k)^{2}+n^{2}$, $K=k^{2}$, $D=d$. Additionally, La-cross codes have weight-$6$ stabilizers consisting of four nearest-neighbor and two symmetrically separated distant qubits – or stabilizer legs. The length of the stabilizer legs is equal to $k$. Since for La-cross codes the generating classical codes are identical, we can drop the indices and use the classical parameter $k$ as a label for different code instances. In fact, for equal rectangular seeds encoding $k$ bits, the resulting HGP code will encode $k^{2}$ logical qubits, hence the parameter $k$ sets both the long-range connectivity and the encoding rate of the resulting code.
@@
Figure 2: Cumulative logical error probability normalized by the number of QEC rounds for k=2k=2(a), k=3k=3(b) and k=4k=4(c) La-cross codes (solid lines) under a fraction $R_{e}=0.98$ of unbiased erasure errors and $R_{p}=1-R_{e}=0.02$ of Pauli errors. A comparison with surface codes under the same noise model and at equal number of physical and logical qubits (dashed lines of the same color and marker style) is also shown. Codes sharing the same number of physical qubits, NN, and logical qubits, $K$, are denoted with the same color. These results show that La-cross codes have high threshold (black stars) and significantly outperform the surface code in terms of QEC performance below physical error probabilities of $\sim 10^{-2}$, for the code distances we have considered. The insets show how the thresholds of La-cross codes increase by increasing the unbiased erasure fraction for $R_{e}=0.40,0.75,0.90,0.98,1.0$. La-cross code instances shown in these plots are: [[34,4,3]],[[52,4,4]],[[100,4,5]],[[130,4,6]][[34,4,3]],[[52,4,4]],[[100,4,5]],[[130,4,6]] for $k=2$; [[45,9,3]],[[65,9,4]],[[149,9,5]],[[225,9,6]][[45,9,3]],[[65,9,4]],[[149,9,5]],[[225,9,6]] for $k=3$; [[80,16,3]],[[106,16,4]],[[136,16,5]],[[208,16,6]][[80,16,3]],[[106,16,4]],[[136,16,5]],[[208,16,6]] for $k=4$. Error bars on the data are standard deviations associated with the Monte Carlo error correction numerical simulations. For both La-cross and surface code, BP+OSD decoder was used.
@@ III.2 Bivariate Bicycle codes
[…]
@@ III.3 Error correction simulations
In this work we have evaluated the QEC performance of quantum LDPC codes via extensive numerical simulationsusing the Clifford simulator Stim [56]. For Clifford-deformed La-cross codes, open boundary conditions, unrotated array configuration [4, 11], and syndrome extraction scheme depicted in Fig. 1(a) (inset) are assumed [11]. Bivariate Bicycle codes are simulated with periodic boundary conditions, unrotated array configuration and standard syndrome extraction scheme. In both cases, the qubit register is first prepared to state $|0\rangle^{\otimes N}$ and – only for Clifford-deformed La-cross codes – Hadamard (H) rotations are applied on alternating data qubits. We apply CNOT/CZ gates between each data and ancilla qubit and then we measure the ancillas in the proper basis. To ensure fault tolerance against errors on the ancilla qubits [4], for all codes, the process is repeated $D$ times for a total of $D$ stabilizer measurement rounds, $D$ being the code distance (see Appendix A for a discussion about the optimal number of rounds.)
We focus on the effects of two-qubit gate errors, which are usually the most detrimental error mechanism, while we assume single-qubit and idling operations to be perfect. We also set the error strength of state preparation and measurement (SPAM) gates to zero, since we are only interested in evaluating two-qubit gate errors. Additionally, it was found in Ref. [26] that including SPAM errors in the simulations only slightly decreases the surface code threshold, while the general sub-threshold behavior is left unchanged. Very similar considerations can be extended to the quantum LDPC codes considered here as well.
We also observe that the choice of neglecting single-qubit and idling errors is physically motivated in neutral atom qubits. In Ref. [32], the fidelity of single-qubit and idling operations are found to be mainly limited by the finite lifetime of the metastable states (few seconds). Specifically, single-qubit gate infidelities are typically more than one order of magnitude lower than two-qubit gates infidelities [48, 32], while idle errors are negligible [7, 32]. All these observations therefore motivate our choice of setting to zero the error strength of single-qubit and idling operations in our QEC numerical simulations.
We assume that each two-qubit gate can experience either a Pauli error with probability $p_{p}=p(1-R_{e})$ or an erasure with probability $p_{e}=pR_{e}$, $R_{e}$ being the fraction of erasure errors. Pauli errors are randomly drawn from $I,X,Y,Z\}^{\otimes 2}\backslash\{I\otimes I\}$ with probability $p_{p}/15$, corresponding to the standard two-qubit depolarizing noise. Unbiased erasure errors are instead modeled as follows: When a two-qubit gate is erased, both atoms – data and ancilla – are reset to the maximally mixed state $I/2$, that is we draw errors at random from $\{I,X,Y,Z\}^{\otimes 2}$ with probability $1/16$, conditionally on an erasure to have occurred. Experimentally, this is consistent with replacing the erased atoms with fresh ones initialized to the maximally mixed state from an external reservoir, e.g. via movable tweezers [26]. In addition, reloading both data and ancilla atoms involved in the erased gate prevents correlated errors to spread during the computation.
@@
For both noise models, we assume that mid-circuit ground-state measurements can be interleaved after any gate layer of the syndrome extraction circuit, ensuring that the failed CNOT/CZ gates are heralded along with the error locations and that the erased atoms can be reloaded, as we model in the simulations. We observe that, despite the increased time overhead of erasure conversion, the initial assumption of negligible idle errors is still well motivated. That is because experimentally errors are heralded via fast and destructive fluorescence imaging on the order of tens of microseconds, which is a much smaller time scale compared to that of idle errors, on the order of few seconds [32].
Figure 3: Cumulative logical error probability normalized by the number of QEC rounds for k=2k=2(a), k=3k=3(b) and k=4k=4(c) La-cross code (colored lines) under a fraction $R_{e}=0.98$ of biased erasure errors and $R_{p}=1-R_{e}=0.02$ of Pauli errors. A comparison with La-cross codes under $R_{e}=0.98$ of unbiased erasure errors (pale lines) is shown, showing the benefit in the threshold (black stars) and in the logical error probability from a such a noise model. We also show a comparison with surface codes (dashed dark lines) with $R_{e}=0.98$ of biased erasure errors, at equal number of physical and logical qubits (dashed lines of the same color and marker style). Codes sharing the same number of physical qubits, NN, and logical qubits, $K$, are denoted with the same color. La-cross codes start to outperform the surface code in terms of QEC performance below physical error probabilities of $\sim 3\times 10^{-2}$, for the code distances we have considered. La-cross code instances shown in these plots are: [[34,4,3]],[[52,4,4]],[[100,4,5]],[[130,4,6]][[34,4,3]],[[52,4,4]],[[100,4,5]],[[130,4,6]] for $k=2$; [[45,9,3]],[[65,9,4]],[[149,9,5]],[[225,9,6]][[45,9,3]],[[65,9,4]],[[149,9,5]],[[225,9,6]] for $k=3$; [[80,16,3]],[[106,16,4]],[[136,16,5]],[[208,16,6]][[80,16,3]],[[106,16,4]],[[136,16,5]],[[208,16,6]] for $k=4$. Error bars on the data are standard deviations associated with the Monte Carlo QEC simulations. For both La-cross and surface code, BP+OSD decoder was used.
@@
circuit-level,
unbiased erasure
$1.6\%$
($1.1\%^{*}$)
$5.6\%$
($4.6\%^{*}$)
$6.8\%$
($5.4\%^{*}$)
$0\% 4.04.0\%$ $5.6\%$ $0.45\%$ $4.2\%$ $5.9\%$ $0.50\%$ $4.6\%$ $6.2\%$
circuit-level,
biased erasure,
native gates
$1.6\%$
($1.1\%^{*}$)
$10.1\%$
($8.7\%^{*}$)
$11.7\%$
($10.6\%^{*}$)
$0.40\%$ $6.8\%$ $9.1\%$ $0.45\%$ $7.0\%$ $9.4\%$ $0.50\%$ $7.2\%$ $9.6\%$
@@
We perform extensive Monte Carlo numerical simulations to probe the code tolerance to errors and use minimum-sum Belief Propagation with Ordered Statistics Decoding (BP+OSD) [57, 58] to process syndrome information of both quantum LDPC codes and surface codes. For all codes, we have set the number of iterations of Belief Propagation to 4104-10 and used Ordered Statistics Decoding at first order. These parameters have been found to optimally trade between decoder accuracy and time overhead [11]. Moreover, although significantly slower than standard matching-based decoders for the surface code, BP+OSD has proven to provide slightly higher threshold and lower logical failure rate [11], therefore we have chosen to show here decoding plots for the surface code obtained with BP+OSD. We also mention that, while the protocol could significantly benfit form shorter decoding times, the huge time overhead of BP+OSD is not an immediate issue for erasure conversion in quantum memory experiments. In fact, error locations are heralded via fast ground state measurements and fresh atoms are reloaded in place of erased ones. Crucially, this does not require real-time decoding, as the erasure detection process is almost perfect, being only limited by the single atom imaging fidelity [32]. Therefore, for quantum memory experiments, offline decoding via BP+OSD represents a viable strategy. Instead, for deep quantum computation with, e.g., gate teleportation, real-time decoding is required and faster decoding strategies for quantum LDPC codes would then become necessary.
We normalize the data by the number of QEC rounds, $D$, i.e. we evaluate the logical failure probability as $P_{L}=1-(1-p_{L})^{1/D}$, $p_{L}$ being the non-normalized cumulative error probability (that is, the probability that any of the $K$ logical qubit fails), calculated as the ratio between the number of decoder failures and the total number of shots. For surface codes, the logical failure rate is computed as $P^{K}_{L}=1-(1-P_{L})^{K}$, $K$ being the number of logical qubits of the compared quantum LDPC code [11].
@@
We show in Fig. 2 results for k=2k=2(a), k=3k=3(b) and k=4k=4(c) Clifford-deformed La-cross codes under a fraction of $R_{e}=0.98$ unbiased erasure errors. We plot their cumulative logical error probability, $P_{L}$, as a function of the physical error probability, $p$, for different code sizes (solid lines). We compare these codes with surface codes of equal/similar number of physical qubits and equal number of logical qubits under the same noise model (dashed lines with the same color and marker style).
We note that, since the encoding rate is the same for both La-cross and surface codes, $K/N_{LDPC}=K/N_{sc}^{K}$, with $N_{sc}^{K}=K\times N_{sc}$, the two codes share the same data and ancilla qubit overhead. In fact, the number of ancilla qubit is $N_{LDPC}^{anc}=N_{LDPC}-K$ for La-cross codes and $N_{sc}^{K,anc}=K\times(N_{sc}-1)=N_{sc}^{K}-K=N_{LDPC}^{anc}$ for surface codes.
@@
It has been shown in Ref. [11] that La-cross codes under depolarizing noise have circuit-level thresholds $p_{th}\approx 0.4-0.5\%$, while for the surface code $p_{th}\approx 1.6\%$. Here, our results show that, under large unbiased erasure fractions, the thresholds of the two code families are comparable. That is, we find thresholds of $p_{th}\approx 4.0-4.6\%$ for La-cross codes [black stars in Fig. 2(a)-(c)] and $p_{th}\approx 5.6\%$ for the surface code. This corresponds to an approximate improvement of $\times 10$ against $\times 5$, showing that the absolute threshold increase of La-cross codes is larger than that of the surface code.